OpenStudy (anonymous):

Solve 6 + 5z = -12(3 – z). Please show your work.

OpenStudy (anonymous):

6 + 5z = -36 + 12z 6+36 = 12z-5z 42 = 7z z = 42/7 z=6

OpenStudy (anonymous):

tthank u so much!

OpenStudy (anonymous):

you're welcome